# Characterisation of distal actions of automorphisms on the space of one-parameter subgroups of Lie groups

### v2

@@ 1. Introduction
In [42], Shah and Yadav have studied the action of automorphisms of a connected Lie groups $G$ on ${\rm Sub}_{G}$ extensively and have shown in particular that for a large class of $G$, which does not have any compact central subgroup of positive dimension, an automorphism $T$ of $G$ acts distally on ${\rm Sub}^{a}_{G}$, the space of closed abelian subgroups of $G$ if and only if $T$ generates a relatively compact subgroup in ${\rm Aut}(G)$. A similar characterisation also holds for $T$ belonging to the class (NC), i.e. those $T$ for which the closure of the $T$-orbit of any discrete cyclic group does not contain the trivial group $\{e\}$. Note that ${\rm Sub}^{a}_{G}$ is very large for many groups $G$; e.g. it is the same as ${\rm Sub}_{G}$ if $G$ is abelian. It also contains all discrete cyclic subgroups. Shah, with Palit in [31], and with Palit and Prajapati in [32], has studied distal actions of automorphisms of a discrete group $G$ on ${\rm Sub}^{a}_{G}$, where $G$ is either polycyclic or a lattice in a connected Lie group.
@@
Let ${\rm Sub}^{c}_{G}$ denote the set of all closed discrete cyclic subgroups of $G$. It is invariant under the action of ${\rm Aut}(G)$ and so is its closure in ${\rm Sub}_{G}$. In [31] and [32], distal actions of automorphisms of ${\rm Sub}^{c}_{\Gamma}$ have been characterised, where $\Gamma$ is a certain type of discrete group, e.g. a discrete polycyclic group or a lattice in a connected Lie group. In a connected Lie group $G$, any closed one-parameter subgroup $H=\{x_{t}\}_{t\in{\mathbb{R}}}$ is the limit of $\{H_{n}\}_{n\in{\mathbb{N}}}$ as $n\to\infty$ in ${\rm Sub}_{G}$, where $H_{n}=\{x_{m/n}\mid m\in{\mathbb{Z}}\}$, $n\in{\mathbb{N}}$, are closed discrete cyclic groups. Thus ${\rm Sub}^{p}_{G}\subset\overline{{\rm Sub}^{c}_{G}}$ (the closure of ${\rm Sub}^{c}_{G}$ in ${\rm Sub}_{G})$ and the latter is contained in ${\rm Sub}^{a}_{G}$. We should note that $\overline{{\rm Sub}^{c}_{G}}$ is much bigger than ${\rm Sub}^{p}_{G}$ as the former contains all discrete cyclic subgroups (see Lemma 2.2). If $T$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$ then $T\in$ (NC). It is easy to see that Theorems 1.1 (2), 1.2 and 1.3 remain valid if ${\rm Sub}^{p}_{G}$ is replaced by $\overline{{\rm Sub}^{c}_{G}}$ in the respective statements.
The following generalises a part of Theorem 4.1 of [42] and also Theorem 1.1 (2).
@@
Note that ${\rm Sub}^{p}_{G}$ is contained in a larger compact set ${\rm Sub}^{co}_{G}$, which is the closure of the set of all closed connected subgroups of $G$. Hence it is easy to show that Theorems 1.11.3 are all valid if ${\rm Sub}^{p}_{G}$ is replaced by ${\rm Sub}^{co}_{G}$. In Theorems 1.4 and 1.5 and Corollary 1.6, the action of the given automorphism or a subgroup of ${\rm Aut}(G)$ is distal on ${\rm Sub}^{co}_{G}$ if and only if it is so on ${\rm Sub}^{p}_{G}$, under the condition on the group and the action on the central torus.
@@
We now fix some notations. Let $G$ be a connected Lie group. Let $R$ (resp. $N$) denote the radical (resp. nilradical) of $G$, i.e. the largest connected solvable (resp. nilpotent) normal subgroup of $G$. Both these subgroup are characteristic in $G$. Recall that $G$ is said to be semisimple if its radical is trivial. In particular, $G/R$ is semisimple. For a closed subgroup $H$ of $G$, let $H^{0}$ denote the connected component of the identity $e$ in $H$, $[H,H]$ denote the commutator subgroup of $H$. Note that $H^{0}$, $[H,H]$ and its closure $\overline{[H,H]}$ are characteristic in $H$. Let ${\mathfrak{g}}$ denote the Lie algebra of the connected Lie group $G$. There is an exponential map $\exp:{\mathfrak{g}}\to G$ which is a continuous map and its restriction to a small open neighbourhood of 00 in ${\mathfrak{g}}$ is a homeomorphism onto an open neighbourhood of the identity $e$ in $G$. Given $T\in{\rm Aut}(G),$ there exists a corresponding Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp\circ\,{\rm d\,}T=T\circ\,\exp$. The map ${\rm d\,}:{\rm Aut}(G)\rightarrow{\rm Aut}({\mathfrak{g}})$, $T\mapsto{\rm d\,}T$ is an injective continuous homomorphism and its image is closed, and in fact, ${\rm d\,}$ is a homeomorphism onto its image (cf. [9], see also [22]) It is known that ${\rm Aut}({\mathfrak{g}})$ is a closed subgroup of ${\rm GL}({\mathfrak{g}})$, and we can view ${\rm Aut}(G)$ as a closed subgroup of ${\rm GL}({\mathfrak{g}})$. The topology on ${\rm Aut}(G)$, when considered as a subspace topology of ${\rm GL}({\mathfrak{g}})$, coincides with the (modified) compact-open topology (see [44]). We refer the reader to [23] for the basic structure theory of Lie groups.
@@ 2. Properties and structure of ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some basic properties of Chabauty topology on the space ${\rm Sub}_{G}$. We also discuss some elementary results about the structure of ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$.
Let $L$ be a locally compact Hausdorff topological group. For a compact set $K$ and an open set $U$ in $L$, let $U_{1}(K)=\{H\in{\rm Sub}_{L}\mid H\cap K=\emptyset\}$ and $U_{2}(U)=\{H\in{\rm Sub}_{L}\mid H\cap U\neq\emptyset\}$. Here, $\{U_{1}(K)\mid K\mbox{ is compact}\}\cup\,\{U_{2}(U)\mid U\mbox{ is open}\}$ is a sub-basis of Chabauty topology on ${\rm Sub}_{L}$. As mentioned earlier, ${\rm Sub}_{L}$ is compact and Hausdorff, and if $L$ is second countable, then ${\rm Sub}_{L}$ is metrisable (cf.  [6], Lemma E.1.1). In particular ${\rm Sub}_{L}$ is metrisable if $L$ is a connected Lie group. The following lemma is known (see e.g. Proposition E.1.2 in [6])
@@ 2. ${\rm(II)}$
The following lemma is well-known (e.g. see Corollary to Lemma 1.2 in [45]; where for a connected Lie group $G$, the conditions for the convergence of a sequence of closed subgroups are equivalent to those in Lemma 2.1 above).
@@
Let $G$ be a connected Lie group. Suppose $\{H_{n}\}_{n\in{\mathbb{N}}}$ in ${\rm Sub}_{G}$ is such that it converges to $H\in{\rm Sub}_{G}$. If $H$ is discrete, then $H_{n}$ is discrete for all large $n$. In particular, the set of discrete subgroups of $G$ is open in ${\rm Sub}_{G}$.
Note that the action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ is continuous for such a $G$ (cf. [42], Lemma 2.4). In particular, if $T$ generates a relatively compact group in ${\rm Aut}(G)$, then it acts distally on $G$ as well as on ${\rm Sub}_{G}$ (see e.g. [31], Lemma 4.2 for the latter statement).
@@
Now ${\rm Sub}^{p}_{G}$ is the closure (in ${\rm Sub}_{G}$) of the set ${\mathcal{P}}_{1}(G)$, i.e.
${\rm Sub}^{p}_{G}:=\overline{{\mathcal{P}}_{1}(G)},$
which is compact and invariant under the action of ${\rm Aut}(G)$.
Note that if any one-parameter subgroup in a connected Lie group is not closed, then its closure is compact and isomorphic to an $n$ dimensional torus $\mathbb{T}^{n}=(\mathbb{S}^{1})^{n}$ for some $n\in{\mathbb{N}}$ with $n\geq 2$. If it is closed and bounded (compact), then it is isomorphic to $\mathbb{S}^{1}$. If it is unbounded (not relatively compact), then it is closed and isomorphic to ${\mathbb{R}}$.
@@
Let $K$ be a compact connected abelian subgroup of $G$. Then the set $F$ of all finite order elements in $K$ is dense in $K$ and each such element of $F$ is contained in a closed one-parameter subgroup of $K$. Since $K$ is monothetic (i.e. it has a dense cyclic subgroup), we get that $K$ is a limit of a sequence of compact one-parameter subgroups of $K$ in ${\rm Sub}_{K}$. Therefore, $K\in{\rm Sub}^{p}_{G}$. Thus (1) holds.
@@
Let $K$ and $\pi$ be as above. Note that $K\in{\rm Sub}^{p}_{G}$ by Lemma 2.3 If $G=K$, then it is easy to see that $(1-4)$ are trivially satisfied.
@@
If $A\in{\mathcal{P}}_{1}(G)$, then $AK$ is closed and $\pi(A)\in{\mathcal{P}}_{1}(G/K)$. Conversely, suppose $B\in{\mathcal{P}}_{1}(G/K$). It is easy to see that $B_{1}:=\pi^{-1}(B)$ is a closed connected nilpotent Lie subgroup of $G$. There exists a neighbourhood $U$ of $e$ such that every element of $U$ is contained in a one-parameter subgroup and $U\setminus K$ is nonempty. Choose any $x\in U\setminus K$, there exists a one-parameter subgroup $\{x_{t}\}$ with $x=x_{1}$. Then $\pi(\{x_{t}\})=B$ and $B_{1}=\{x_{t}\}K$. As $K$ is central in $G$ we get that $B_{1}$ is, in fact, abelian. If $\{x_{t}\}$ is closed, then it belongs to ${\mathcal{P}}_{1}(G)$ and $\pi(\{x_{t}\})=B$. Now suppose it is not closed, then its closure is compact, and hence $B_{1}$ is compact. This implies that the set of torsion elements is dense in $B_{1}$. Since $U\setminus K$ is open, we can choose $y\in U\setminus K$ such that $y^{n}=e$, for some $n\in{\mathbb{N}}\setminus\{1\}$. Let $\{y_{t}\}$ be a one-parameter subgroup in $B_{1}$ such that $y_{1}=y$. Then $\pi(\{y_{t}\})=B$. Here, $\{y_{t}\}$ is closed as $y_{1}^{n}=e$. Therefore, ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$, i.e. (2) holds.
Let $H\in{\rm Sub}^{p}_{G}$. Then $A_{n}\to H$ for some $\{A_{n}\}\subset{\mathcal{P}}_{1}(G)$. By (2), we have that $\{\pi(A_{n})\}\subset{\mathcal{P}}_{1}(G/K)$, and as $K$ is compact, $\pi(A_{n})\to\pi(H)$. Therefore, $\pi(H)\in{\rm Sub}^{p}_{G/K}$. Conversely, let $B\in{\rm Sub}^{p}_{G/K}$. There exist $B_{n}\in{\mathcal{P}}_{1}(G/K)$, $n\in{\mathbb{N}}$, such that $B_{n}\to B$. By (2), there exist $A_{n}\in{\mathcal{P}}_{1}(G)$ such that $\pi(A_{n})=B_{n}$, $n\in{\mathbb{N}}$. Let $A$ be a limit point of $\{A_{n}\}$ in ${\rm Sub}_{G}$. Then $A\in{\rm Sub}^{p}_{G}$ and $B=\pi(A)$. Thus (3) holds.
Let ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$. It is easy to show that ${\mathcal{S}}(K)$ is closed in ${\rm Sub}^{p}_{G}$, and hence it is compact. Observe that the set ${\mathcal{S}}(K)$ is same as the set $\{HK\mid H\in{\rm Sub}^{p}_{G}\}$. For $H_{1},H_{2}\in{\rm Sub}_{G}$, if $\pi(H_{1})=\pi(H_{2})$, then $H_{1}K=H_{2}K$. Therefore, the map $H\mapsto\pi(H)$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is injective, and it easily follows from (3) that it is surjective. Also if $H_{n}\to H$ in ${\rm Sub}_{G}$, then $\pi(H_{n})\to\pi(H)$ in ${\rm Sub}_{G/K}$. Since ${\mathcal{S}}(K)$ is compact and ${\rm Sub}^{p}_{G/K}$ is Hausdorff, we have that the map induced by $\pi$ from ${\mathcal{S}}(K)$ to ${\rm Sub}^{p}_{G/K}$ is a homeomorphism. Thus (4) holds. ∎

### v3

@@ 1. Introduction
In [42], Shah and Yadav have studied the action of automorphisms of a connected Lie group $G$ on ${\rm Sub}_{G}$ extensively and have shown in particular that for a large class of $G$, which does not have any compact central subgroup of positive dimension, an automorphism $T$ of $G$ acts distally on ${\rm Sub}^{a}_{G}$, the space of closed abelian subgroups of $G$, if and only if $T$ generates a relatively compact subgroup in ${\rm Aut}(G)$. A similar characterisation also holds for $T$ belonging to the class (NC), i.e. those $T$ for which the closure of the $T$-orbit of any discrete cyclic group does not contain the trivial group $\{e\}$. Note that ${\rm Sub}^{a}_{G}$ is very large for many groups $G$; e.g. it is the same as ${\rm Sub}_{G}$ if $G$ is abelian. It also contains all discrete cyclic subgroups. Shah, with Palit in [31], and with Palit and Prajapati in [32], has studied distal actions of automorphisms of a discrete group $G$ on ${\rm Sub}^{a}_{G}$, where $G$ is either polycyclic or a lattice in a connected Lie group.
@@
Let ${\rm Sub}^{c}_{G}$ denote the set of all discrete cyclic subgroups of $G$. It is invariant under the action of ${\rm Aut}(G)$ and so is its closure in ${\rm Sub}_{G}$. In [31] and [32], distal actions of automorphisms of ${\rm Sub}^{c}_{\Gamma}$ have been characterised, where $\Gamma$ is a certain type of discrete group, e.g. a discrete polycyclic group or a lattice in a connected Lie group. In a connected Lie group $G$, any closed one-parameter subgroup $H=\{x_{t}\}_{t\in{\mathbb{R}}}$ is the limit of $\{H_{n}\}_{n\in{\mathbb{N}}}$ as $n\to\infty$ in ${\rm Sub}_{G}$, where $H_{n}=\{x_{m/n}\mid m\in{\mathbb{Z}}\}$, $n\in{\mathbb{N}}$, are discrete cyclic groups. Thus ${\rm Sub}^{p}_{G}\subset\overline{{\rm Sub}^{c}_{G}}$ (the closure of ${\rm Sub}^{c}_{G}$ in ${\rm Sub}_{G})$ and the latter is contained in ${\rm Sub}^{a}_{G}$. We should note that $\overline{{\rm Sub}^{c}_{G}}$ is much bigger than ${\rm Sub}^{p}_{G}$ as the former contains all discrete cyclic subgroups (see Lemma 2.2). If $T$ acts distally on $\overline{{\rm Sub}^{c}_{G}}$, then $T\in$ (NC). It is easy to see that Theorems 1.1 (2), 1.2 and 1.3 remain valid if ${\rm Sub}^{p}_{G}$ is replaced by $\overline{{\rm Sub}^{c}_{G}}$ in the respective statements.
The following generalises a part of Theorem 4.1 of [42] and also extends Theorem 1.1 (2).
@@
Note that ${\rm Sub}^{p}_{G}$ is contained in a larger compact set ${\rm Sub}^{co}_{G}$, which is the closure of the set of all closed connected subgroups of $G$. Hence it is easy to show that Theorems 1.11.3 are all valid if ${\rm Sub}^{p}_{G}$ is replaced by ${\rm Sub}^{co}_{G}$. In Theorems 1.4 and 1.5 and Corollary 1.6, the action of a given automorphism or a subgroup of ${\rm Aut}(G)$ is distal on ${\rm Sub}^{co}_{G}$ if and only if it is so on ${\rm Sub}^{p}_{G}$, under the condition on the group and the action on the central torus.
@@
We now fix some notations. Let $G$ be a connected Lie group. Let $R$ (resp. $N$) denote the radical (resp. nilradical) of $G$, i.e. the largest connected solvable (resp. nilpotent) normal subgroup of $G$. Both these subgroup are characteristic in $G$. Recall that $G$ is said to be semisimple if its radical is trivial. In particular, $G/R$ is semisimple unless $G$ is solvable. For a closed subgroup $H$ of $G$, let $H^{0}$ denote the connected component of the identity $e$ in $H$, $[H,H]$ denote the commutator subgroup of $H$. Note that $H^{0}$, $[H,H]$ and its closure $\overline{[H,H]}$ are characteristic in $H$. Let ${\mathfrak{g}}$ denote the Lie algebra of the connected Lie group $G$. There is an exponential map $\exp:{\mathfrak{g}}\to G$ which is a continuous map and its restriction to a small open neighbourhood of 00 in ${\mathfrak{g}}$ is a homeomorphism onto an open neighbourhood of the identity $e$ in $G$. Given $T\in{\rm Aut}(G),$ there exists a corresponding Lie algebra automorphism ${\rm d\,}T:{\mathfrak{g}}\to{\mathfrak{g}}$ such that $\exp\circ\,{\rm d\,}T=T\circ\,\exp$. The map ${\rm d\,}:{\rm Aut}(G)\rightarrow{\rm Aut}({\mathfrak{g}})$, $T\mapsto{\rm d\,}T$ is an injective continuous homomorphism and its image is closed, and in fact, ${\rm d\,}$ is a homeomorphism onto its image (cf. [9], see also [22]) It is known that ${\rm Aut}({\mathfrak{g}})$ is a closed subgroup of ${\rm GL}({\mathfrak{g}})$, and we can view ${\rm Aut}(G)$ as a closed subgroup of ${\rm GL}({\mathfrak{g}})$. The topology on ${\rm Aut}(G)$, when considered as a subspace topology of ${\rm GL}({\mathfrak{g}})$, coincides with the (modified) compact-open topology (see [44]). We refer the reader to [23] for the basic structure theory of Lie groups.
@@ 2. Properties and structure of ${\mathrm{Sub}}_{G}$ and ${\mathrm{Sub}}^{p}_{G}$
In this section, we discuss some basic properties of the Chabauty topology on the space ${\rm Sub}_{G}$. We also discuss some elementary results about the structure of ${\rm Sub}_{G}$ and ${\rm Sub}^{p}_{G}$.
Let $L$ be a locally compact Hausdorff topological group. For a compact set $K$ and an open set $U$ in $L$, let $U_{1}(K)=\{H\in{\rm Sub}_{L}\mid H\cap K=\emptyset\}$ and $U_{2}(U)=\{H\in{\rm Sub}_{L}\mid H\cap U\neq\emptyset\}$. Here, $\{U_{1}(K)\mid K\mbox{ is compact}\}\cup\,\{U_{2}(U)\mid U\mbox{ is open}\}$ is a sub-basis of the Chabauty topology on ${\rm Sub}_{L}$. As mentioned earlier, ${\rm Sub}_{L}$ is compact and Hausdorff, and if $L$ is second countable, then ${\rm Sub}_{L}$ is metrisable (cf.  [6], Lemma E.1.1). In particular ${\rm Sub}_{L}$ is metrisable if $L$ is a connected Lie group. For a closed subgroup $H$ of $L$, it is easy to see that ${\rm Sub}_{H}$ carries the subspace topology of ${\rm Sub}_{L}$. The following lemma is known (see e.g. Proposition E.1.2 in [6])
@@ 2. ${\rm(II)}$
The following lemma is a consequence of the well-known result that the set of discrete subgroups of a connected Lie group $G$ is open in ${\rm Sub}_{G}$. We will give a short proof for the sake of completeness.
@@
Let $G$ be a connected Lie group. Then the trivial subgroup $\{e\}$ is isolated in ${\rm Sub}^{co}_{G}$, and so in ${\rm Sub}^{p}_{G}$. Moreover, ${\rm Sub}^{co}_{G}$ does not contain any non-trivial discrete subgroup.
Proof.
Let $\mathfrak{C}$ be the set of closed connected subgroups of $G$. We have ¯$\overline{\mathfrak{C}}={\rm Sub}^{co}_{G}$. By Proposition E.1.5 in [6] (see also Proposition 3.4(iii) in [8], or also Corollary of Lemma 1.2 in [45]) the set $\mathfrak{D}$ of discrete subgroups of $G$ is open in ${\rm Sub}_{G}$ and so $\{\{e\}\}\subseteq\mathfrak{D}\cap{\rm Sub}^{co}_{G}=\mathfrak{D}\cap\overline{\mathfrak{C}}\subseteq\overline{\mathfrak{D}\cap\mathfrak{C}}=\{\{e\}\}$. Hence the assertions hold. ∎
Note that the action of ${\rm Aut}(G)$ on ${\rm Sub}_{G}$ is continuous for connected Lie groups $G$ (cf. [42], Lemma 2.4). In particular, if $T$ generates a relatively compact group in ${\rm Aut}(G)$, then it acts distally on $G$ as well as on ${\rm Sub}_{G}$ (see e.g. [31], Lemma 4.2 for the latter statement).
@@
Now ${\rm Sub}^{p}_{G}$ is the closure (in ${\rm Sub}_{G}$) of the set ${\mathcal{P}}_{1}(G)$, i.e. ${\rm Sub}^{p}_{G}:=\overline{{\mathcal{P}}_{1}(G)}$, which is compact and invariant under the action of ${\rm Aut}(G)$.
Note that if any one-parameter subgroup in a connected Lie group is not closed, then its closure is compact and isomorphic to an $n$-dimensional torus $\mathbb{T}^{n}=(\mathbb{S}^{1})^{n}$ for some $n\in{\mathbb{N}}$ with $n\geq 2$. If it is closed and bounded (compact), then it is isomorphic to $\mathbb{S}^{1}$. If it is unbounded (not relatively compact), then it is closed and isomorphic to ${\mathbb{R}}$.
@@
Let $K$ be a compact connected abelian subgroup of $G$. Then the set $F$ of all finite order elements in $K$ is dense in $K$ and each such element of $F$ is contained in a closed one-parameter subgroup of $K$. Since $K$ is monothetic (i.e. it has a dense cyclic subgroup), we get that $K$ is a limit of a sequence of compact one-parameter subgroups of $K$ in ${\rm Sub}_{K}$. Therefore, $K\in{\rm Sub}^{p}_{K}\subset{\rm Sub}^{p}_{G}$. Thus (1) holds.
@@
Let $K$ and $\pi$ be as above. Note that $K\in{\rm Sub}^{p}_{G}$ by Lemma 2.3. If $G=K$, then it is easy to see that $(1-4)$ are trivially satisfied.
@@
If $A\in{\mathcal{P}}_{1}(G)$, then $AK$ is closed and $\pi(A)\in{\mathcal{P}}_{1}(G/K)$. Conversely, suppose $B\in{\mathcal{P}}_{1}(G/K$) is nontrivial. It is easy to see that $B_{1}:=\pi^{-1}(B)$ is a closed connected nilpotent Lie subgroup of $G$. There exists a neighbourhood $U$ of $e$ such that every element of $U$ is contained in a one-parameter subgroup and $U\setminus K$ is nonempty. Choose any $x\in U\setminus K$, there exists a one-parameter subgroup $\{x_{t}\}$ with $x_{1}=x$. Then $\pi(\{x_{t}\})=B$ and $B_{1}=\{x_{t}\}K$. As $K$ is central in $G$ we get that $B_{1}$ is, in fact, abelian. If $\{x_{t}\}$ is closed, then it belongs to ${\mathcal{P}}_{1}(G)$ and $\pi(\{x_{t}\})=B$. Now suppose it is not closed, then its closure is compact, and hence $B_{1}$ is compact. This implies that the set of torsion elements is dense in $B_{1}$. Since $U\setminus K$ is open, we can choose $y\in B_{1}\cap(U\setminus K)$ such that $y^{n}=e$, for some $n\in{\mathbb{N}}\setminus\{1\}$. Let $\{y_{t}\}$ be a one-parameter subgroup in $B_{1}$ such that $y_{1}=y$. Then $\pi(\{y_{t}\})=B$. Here, $\{y_{t}\}$ is closed as $y_{1}^{n}=e$. Therefore, ${\mathcal{P}}_{1}(G/K)=\{\pi(A)\mid A\in{\mathcal{P}}_{1}(G)\}$, i.e. (2) holds.
Let $\pi^{\prime}:{\rm Sub}^{p}_{G}\to{\rm Sub}^{p}_{G/K}$ be the map induced by $\pi$. Then $\pi^{\prime}$ is a continuous closed map. From (2), we get that ${\rm Sub}^{p}_{G/K}=\overline{{\mathcal{P}}_{1}(G/K)}=\pi^{\prime}(\overline{{\mathcal{P}}_{1}(G)})=\pi^{\prime}({\rm Sub}^{p}_{G})=\{\pi(H)\mid H\in{\rm Sub}^{p}_{G}\}$. Thus (3) holds.
Let ${\mathcal{S}}(K):=\{HK\mid H\in{\rm Sub}^{p}_{G}\}$. By (1), ${\mathcal{S}}(K)$ is closed in ${\rm Sub}^{p}_{G}$, and hence it is compact. Observe that ${\mathcal{S}}(K)=\{H\in{\rm Sub}^{p}_{G}\mid HK=KH=H\}$. By (3), $\pi^{\prime}({\mathcal{S}}(K))=\pi^{\prime}({\rm Sub}^{p}_{G})={\rm Sub}^{p}_{G/K}$ and it is obvious that $\pi^{\prime}|_{{\mathcal{S}}(K)}$ is continuous and bijective, and hence it is a homeomorphism. Thus (4) holds. ∎